Product-Costing Accuracy, Plantwide and Departmental Rates, ABC

Escuha Company produces two type of calculators: scientific and business. Both products pass through two producing departments. The business calculator is by far the most popular. The following data have been gathered for these two products:

Product-Related Data
Scientific Business
Units produced per year 77,000 770,000
Prime costs $265,000 $2,650,000
Direct labor hours 101,000 1,010,000
Machine hours 49,000 490,000
Production runs 100 150
Inspection hours 2,000 3,000
Maintenance hours 2,250 9,000
Department Data
Department 1 Department 2
Direct labor hours:
  Scientific calculator 76,700 24,300
Business calculator 121,200 888,800
Total 197,900 913,100
Machine hours:
Scientific calculator 24,600 24,400
Business calculator 408,360 81,640
Total 432,960 106,040
Overhead costs:
Setup costs $230,000 $230,000
Inspection costs 170,000 170,000
Power 236,000 155,000
Maintenance 191,000 247,750
Total $827,000 $802,750

Required:

1. Compute the overhead cost per unit for each product using a plantwide, unit-based rate using direct labor hours. Round your calculations and answers to the nearest cent.

Plantwide rate $ per direct labor hour
Overhead cost per unit:
Scientific $ per unit
Business $ per unit

2. Compute the overhead cost per unit for each product using departmental rates. In calculating departmental rates, use machine hours for Department 1 and direct labor hours for Department 2. Round departmental rates and the overhead cost per unit to the nearest cent if rounding is required.

Departmental rates:
Department 1 $ per machine hour
Department 2 $ per direct labor hour
Overhead cost per unit:
Scientific $ per unit
Business $ per unit

Repeat using direct labor hours for Department 1 and machine hours for Department 2. Round departmental rates and the overhead cost per unit to the nearest cent if rounding is required.

Departmental rates:
Department 1 $ per direct labor hour
Department 2 $ per machine hour
Overhead cost per unit:
Scientific $ per unit
Business $ per unit

3. Compute the overhead cost per unit for each product using activity-based costing. Round your calculations and answers to the nearest cent.

Overhead Cost
Scientific $ per unit
Business $ per unit
Answer #1

Solution

Escuha Company

  1. Computation of the overhead cost per unit for each product using a plantwide, unit-based rate using direct labor hours:

Plantwide rate

$ per direct labor hour

Overhead cost per unit:

Scientific

$1.92 per unit

Business

$1.92 per unit

Computations:

Plantwide overhead rate = total overhead costs/direct labor hours

Total overhead cost = department 1 costs + department 2 costs

= $827,000 + 802,750 = $1,629,750

Total direct labor hours = department 1 DLH + department 2 DLH

= 197,900 + 913,100 = 1,111,000

Plandwide overhead rate = $1,629,750/1,111,000 = $1.47 per DLH

Overhead cost assigned = plantwide overhead rate x DLH used

Scientific Calculator –

Direct labor hours = 76,700 + 24,300 = 101,000

Overhead cost assigned = $1.47 x 101,000 = $148,160

Number of units = 77,000

Overhead cost per unit = $148,160/77,000 = $1.92

Business Calculator –

Direct labor hours = 121,200 + 888,800 = 1,010,000

Overhead cost assigned = $1.47 x 1,010,000 = $1,481,590

Number of units = 770,000

Overhead cost per unit = $1,481,590/770,000 = $1.92

  1. Computation of the overhead cost per unit for each product using a plantwide, unit based rate using department rates:

Department 1 – Machine hours; Department 2 – Direct labor hours

Department Rates

Department 1

$1.91 per MH

Department 2

$0.88 per DLH

Overhead Cost per Unit

Scientific

$0.89

Business

$2.02

Computations:

Department 1-

Total overhead costs = $827,000

Total machine hours = 432,960

Overhead rate per MH = $827,000/432,960 = $1.91

Department 1 overhead cost assigned to

Scientific Calculator –

MH = 24,600

Overhead rate = $1.91

Assigned overhead cost = $1.91 x 24,600 = $46,989

Number of units = 77,000

Department 1 overhead cost per unit = $0.61

Business Calculator –

MH = 408,360

Overhead rate = $1.91

Assigned overhead cost = $1.91 x 408,360 = $780,011

Number of units = 770,000

Department 1 overhead cost per unit = 780,011/770,000 = $1.01

Department 2-

Total overhead costs = $802,750

Total direct labor hours = 913,100

Overhead rate per DLH = $802,750/913,100 = $0.88

Department 2 overhead cost assigned to

Scientific Calculator –

DLH = 24,300

Overhead rate = $0.88

Assigned overhead cost = $0.88 x 24,300 = $21,363

Number of units = 77,000

Department 2 overhead cost per unit = $0.28

Business Calculator –

DLH = 888,800

Overhead rate = $0.88

Assigned overhead cost = $0.88 x 888,800 = $781,387

Number of units = 770,000

Department 2 overhead cost per unit = 781,387/770,000 = $1.015

Total overhead cost per unit –

Scientific Calculator -
department 1 = $0.61

Department 2 = $0.28

Total overhead cost per unit = $0.89

Business Calculator –

Department 1 = $1.01

Department 2 = $1.015

Total overhead cost per unit = $2.02

Computation of the overhead cost per unit for each product using a plantwide, unit-based rate using department rates:

Department 1 – Direct labor hours;          Department 2 – Machine hours

Department Rates

Department 1

$4.18 per DLH

Department 2

$7.57 per MH

Overhead Cost per Unit

Scientific

$6.56

Business

$1.46

Computations:

Department 1-

Total overhead costs = $827,000

Total direct labor hours = 197,900

Overhead rate per MH = $827,000/197,900 = $4.18

Department 1 overhead cost assigned to

Scientific Calculator –

DLH = 76,700

Overhead rate = $4.18

Assigned overhead cost = $4.18 x 76,700 = $320,520

Number of units = 77,000

Department 1 overhead cost per unit = 320,520/77,000 = $4.16

Business Calculator –

DLH = 121,200

Overhead rate = $4.18

Assigned overhead cost = $4.18 x 121,200 = $506,480

Number of units = 770,000

Department 1 overhead cost per unit = 506,480/770,000 = $0.66

Department 2-

Total overhead costs = $802,750

Total Machine hours = 106,040

Overhead rate per MH = $802,750/106,040 = $7.57

Department 2 overhead cost assigned to

Scientific Calculator –

MH = 24,400

Overhead rate = $7.57 per MH

Assigned overhead cost = $7.57 x 24,400 = $184,714

Number of units = 77,000

Department 2 overhead cost per unit = $2.40

Business Calculator –

MH = 81,640

Overhead rate = $7.57 per MH

Assigned overhead cost = $7.57 x 81,640 = $618,036

Number of units = 770,000

Department 2 overhead cost per unit = 618,036/770,000 = $0.80

Total overhead cost per unit –

Scientific Calculator -
department 1 = $4.16

Department 2 = $2.40

Total overhead cost per unit = $6.56

Business Calculator –

Department 1 = $0.66

Department 2 = $0.80

Total overhead cost per unit = $1.46

  1. Computation of the overhead cost per unit for each product using activity – based costing:

Overhead cost per unit

Scientific

$5.76

Business

$1.54

Computations:

Activity

Cost

Activity Usage

Activity rate = cost/usage

Setup

$460,000

250 production runs

$1,840 per production hour

Inspection

$340,000

5,000 inspection hours

$68 per inspection hour

Power

$391,000

539,000 machine hours

$0.73 per machine hour

Maintenance

$438,750

11,250 maintenance hours

$39 per maintenance hour

Scientific Calculator

Activity

Activity Rate

Activity Usage

Cost Assigned

Setup

$1,840 per production run

100 runs

$184,000

Inspection

$68 per inspection hour

2,000 inspection hours

$136,000

Power

$0.73 per MH

49,000 MH

$35,770

Maintenance

$39 per maintenance hour

2,250 maintenance hours

$87,750

Total

$443,520

number of units

77,000

Overhead cost per unit

$5.76

Business Calculator

Activity

Activity Rate

Activity Usage

Cost Assigned

Setup

$1,840 per production run

150 runs

$276,000

Inspection

$68 per inspection hour

3,000 inspection hours

$204,000

Power

$0.73 per MH

490,000 MH

$357,700

Maintenance

$39 per maintenance hour

9,000 maintenance hours

$351,000

Total

$1,188,700

number of units

770,000

Overhead cost per unit

$1.54